# Contextualized Messages Boost Graph Representations

### v2

@@ Abstract
Graph neural networks (GNNs) have gained significant attention in recent years for their ability to process data that may be represented as graphs. This success has prompted several studies to explore the representational capability of GNNs based on the graph isomorphism task. These works inherently assume a countable node feature representation, potentially limiting their applicability. Interestingly, only a few theoretical works study GNNs with uncountable node feature representation. This paper presents a novel perspective on the representational capability of GNNs across all levels – node-level, neighborhood-level, and graph-level – when the space of node feature representation is uncountable. Specifically, it relaxes the injective requirement in previous works by employing an implicit pseudometric distance on the space of input to create a soft-injective function. This allows distinct inputs to produce similar outputs only if the pseudometric deems the inputs to be sufficiently similar on some representation, which is often useful in practice. As a consequence, a novel soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes non-linear and contextualized transformation of neighborhood feature representations is proposed. A mathematical discussion on the relationship between SIR-GCN and widely used GNNs is then laid out to put the contribution in context, establishing SIR-GCN as a generalization of classical GNN methodologies. Experiments on synthetic and benchmark datasets demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
@@ 1 Introduction
Graph neural networks (GNNs) constitute a class of deep learning models designed to process data that may be represented as graphs. These models are well-suited for node, edge, and graph property prediction tasks across various domains including social networks, molecular graphs, and biological networks, among others [12, 19]. GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation [14, 47, 48]. This allows the model to encapsulate both the network structure and the broader node contexts. Moreover, a graph readout function is employed to pool the individual node feature representation and create a representation for the entire graph [27, 32, 47, 50].
@@
Among the most widely used GNNs in literature include the graph convolution network (GCN) [25], graph sample and aggregate (GraphSAGE) [15], graph attention network (GAT) [7, 40], and graph isomorphism network (GIN) [47] which largely fall under the message-passing neural networks (MPNNs) [14] framework. These models have gained popularity due to their simplicity and remarkable performance across various applications [12, 17, 19, 22, 24, 28]. Improvements are also constantly being proposed to achieve state-of-the-art performance [4, 6, 21, 31, 38, 44, 49].
@@
Notably, these advances are mainly driven by heuristics and empirical results. Nonetheless, several studies have also begun exploring the representational capability of GNNs [2, 4, 5, 11, 13, 34]. Most of these works analyzed GNNs in relation to the graph isomorphism task. Xu et al. [47] was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test [46]. Subsequent works build upon their results by considering extensions to the original 1-WL test. However, these results only hold with countable node feature representation which potentially limits their applicability. Meanwhile, Corso et al. [11] proposed using multiple aggregators to create powerful GNNs when the space of node feature representation is uncountable. Interestingly, there has been no significant theoretical progress since this work.
@@
This paper presents a simple yet novel perspective on the representational capability of GNNs when the space of node feature representation is uncountable. The key idea is to define an implicit pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs only if the distance between the inputs is sufficiently small on some representation. This idea is explored across all levels – node-level, neighborhood-level, and graph-level. Based on the results, a novel soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes the non-linear and contextualized transformation of neighborhood feature representations is proposed. The mathematical relationship between SIR-GCN and popular GNNs in literature is also presented to underscore the advantages of the proposed model. Experiments on synthetic and benchmark datasets in node and graph property prediction tasks then highlight the expressivity of SIR-GCN in closing.
@@ 2 Graph neural networks
Let $\mathcal{G}=\left(\mathcal{N}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{N}_{\mathcal{G}}$ the set of nodes adjacent to node $u\in\mathcal{N}_{\mathcal{G}}$. The subscript $\mathcal{G}$ will be omitted whenever the context is clear. Suppose $\mathcal{H}$ is the space of node feature representation, henceforth feature, and $\bm{h_{u}}\in\mathcal{H}$ is the feature of node $u$. A GNN following the message-passing scheme can be expressed mathematically as
@@
where AGG and COMB are some aggregation and combination strategies, respectively, $\bm{H_{u}}$ is the multiset [47] of neighborhood features for node $u$, $\bm{a_{u}}$ is the aggregated neighborhood features for node $u$, and $\bm{h^{*}_{u}}$ is the updated feature for node $u$. Since AGG takes arbitrary-sized multisets of neighborhood features as input and transforms them into a single feature, it may be considered a hash function. Hence, aggregation and hash functions shall be used interchangeably throughout the paper.
@@
When $\mathcal{H}$ is countable, Xu et al. [47] showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
@@
is injective or unique for each multiset of neighborhood features $\bm{H}$ of bounded size. This result forms the theoretical basis of GIN.
Meanwhile, the result above no longer holds when $\mathcal{H}$ is uncountable. In this setting, Corso et al. [11] proved that if $\bigoplus$ comprises multiple aggregators (e.g. mean, standard deviation, max, and min), the hash function
@@
produces a unique output for every $\bm{H}$ of bounded size. This finding laid the foundation for the principal neighborhood aggregation (PNA) [11]. Notably, for this result to hold, the number of aggregators in $\bigoplus$ must also scale with the size of the multiset of neighborhood features $\bm{H}$, which may be infeasible for large and dense graphs.
@@ 3 Soft-injective functions
Theorem 1 presents an alternative to injective functions when the space of node features $\mathcal{H}$ is uncountable. It considers a soft relaxation to injectivity, henceforth soft-injectivity, which is often useful in practice, especially in classification tasks.
@@
Let $\mathcal{H}$ be a non-empty set with a pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$, there exists $\delta_{1}>\delta_{2}>0$ satisfying
$\delta_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\delta_{1}\implies\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (4)
Theorem 1 shows that, for every node $u$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function operating on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ only if $d_{u}$ determines $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}$ to be sufficiently similar on some representation. The lower bounds $\delta_{2}$ and $\varepsilon_{2}$ assert the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bounds $\delta_{1}$ and $\varepsilon_{1}$ ensure $g_{u}$ maintains the relationship of elements of $\mathcal{H}$ with respect to $d_{u}$. The feature map $g_{u}$ is then said to be soft-injective. An illustration is provided in Fig. 1.
$\delta$$g_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}}\right)$$g_{u}\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)$$g_{u}\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
(a) Embedded feature space $\mathcal{S}$.
@@
(b) Input feature space $\mathcal{H}$.
Figure 1: Soft-injective feature map $g_{u}$ with pseudometric $d_{u}$.
It is worth noting that if $d_{u}$ is squared Euclidean, then Theorem 1 becomes trivial. Nevertheless, the result becomes non-trivial for other choices of $d_{u}$. Corollary 1 extends this result for multisets.
@@
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}}),$ (5)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}$ and bounded, equinumerous multisets of $\mathcal{H}$ $\bm{H}^{(1)},\bm{H}^{(2)}$. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{H}^{(1)},\bm{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$, there exists $\delta_{1}>\delta_{2}>0$ satisfying
$\delta_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\delta_{1}\implies\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1},$ (6)
@@
3.1 Soft-isomorphic relational graph convolution network
Corollary 1 shows that, for every node $u$, given a pseudometric distance $d_{u}$ on $\mathcal{H}$ with a corresponding pseudometric distance $D_{u}$ on multisets of $\mathcal{H}$ defined in Eqn. 5, there exists a corresponding feature map $g_{u}$ and soft-injective hash function $G_{u}$ defined in Eqn. 7 that produces similar outputs for distinct multisets of neighborhood features $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ only if $D_{u}$ deems $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ to be sufficiently similar on some representation. Similarly, the lower and upper bounds guarantee the ability of $G_{u}$ to separate equinumerous multisets of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while maintaining the relationship with respect to $D_{u}$. In this setting, the feature map $g_{u}$ may be interpreted as the message function [14] of the aggregation strategy that transforms the individual neighborhood features. Meanwhile, the distance function $D_{u}$ may be interpreted as a kernel distance [23].
Definition 1 (Collision).
Let $G$ be a function. If $G\left(\bm{H}^{(1)}\right)=G\left(\bm{H}^{(2)}\right)$ and $\bm{H}^{(1)}\neq\bm{H}^{(2)}$, a collision is said to have occurred.
To illustrate the utility of $D_{u}$, suppose node $u$ has two neighbors $v_{1}$, $v_{2}$. If $d_{u}$ is the squared Euclidean distance, then a corresponding message function $g_{u}$ is linear. Fig. 2(a) presents the contour plot of the corresponding hash function $G_{u}$, highlighting potential issues arising from hash collisions. Specifically, consider two multisets $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$. If $\bm{h_{v_{1}}}^{(1)}+\bm{h_{v_{2}}}^{(1)}=\bm{h_{v_{1}}}^{(2)}+\bm{h_{v_{2}}}^{(2)}$, then a hash collision occurs and $G_{u}$ produces identical aggregated neighborhood features even if $\bm{H_{u}}^{(1)}$ and $\bm{H_{u}}^{(2)}$ are fundamentally dissimilar on some representation for a given task.
@@ 3.1 Soft-isomorphic relational graph convolution network
In general, hash collisions may occur when aggregating neighborhood features due to $\mathcal{H}$ being uncountable. Within the current framework, one may simply encode knowledge about node features into the pseudometric $D_{u}$. As a result, only the regions determined by $D_{u}$ to be similar may produce similar aggregated neighborhood features, making collisions more informative and controlled. This also corresponds to a more complex and non-linear message function $g_{u}$. To illustrate, if node features represent a zero-mean score, $d_{u}$ may be defined as the squared Euclidean distance of the squared score. A corresponding hash function $G_{u}$ in Fig. 2(b) may then be used to detect potentially anomalous neighborhoods since hash collisions are meaningful for this task.
@@
It is worth noting, however, that Corollary 1 holds for every node $u\in\mathcal{N}$ independently. Hence, different nodes may correspond to different $D_{u}$ and $G_{u}$. For simplicity, one may consider only a single pseudometric compactly defined as
$D^{2}\left(\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)};\bm{h_{u}}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(1)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(2)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}d(\bm{h},\bm{h^{\prime}};\bm{h_{u}})$ (8)
with a corresponding soft-injective hash function
$G\left(\bm{H_{u}};\bm{h_{u}}\right)=\sum_{\bm{h}\in\bm{H_{u}}}g\left(\bm{h};\bm{h_{u}}\right)$ (9)
@@
for every node $u\in\mathcal{N}$. This approach preserves the interpretation of $G$ as an aggregation or hash function with an underlying pseudometric distance $D$ that guides and controls hash collisions. The integration of $\bm{h_{u}}$ also allows for the interpretation of $g$ as a relational message function guiding how features of the key (neighboring) nodes $\bm{h}\in\bm{H_{u}}$ are to be embedded and transformed based on the features of the query (center) node $\bm{h_{u}}$. This provides additional context to hash collisions and makes the message function and pseudometric anisotropic [12] or adaptive with respect to the query node. Figs. 2(c) and 2(d) demonstrate this idea where the introduction of a bias term, assuming a function of $\bm{h_{u}}$, shifts the contour plot and produces different aggregated neighborhood features for identical neighborhood features. Nevertheless, one may opt to inject stochasticity into the node features to distinguish between nodes with identical features and neighborhood features and imitate having distinct hash functions $G_{u}$ for every node $u$ with high probability [34].

### v3

# No License

@@ Abstract
Graph neural networks (GNNs) have gained significant attention in recent years for their ability to process data that may be represented as graphs. This has prompted several studies to explore their representational capability based on the graph isomorphism task. These works inherently assume a countable node feature representation, potentially limiting their applicability. Interestingly, only a few study GNNs with uncountable node feature representation. In the paper, a novel perspective on the representational capability of GNNs is investigated across all levels—node-level, neighborhood-level, and graph-level—when the space of node feature representation is uncountable. More specifically, the strict injective and metric requirements are softly relaxed by employing a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the pseudometric deems the inputs to be sufficiently similar on some representation. As a consequence, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) that emphasizes the contextualized transformation of neighborhood feature representations via anisotropic and dynamic message functions is proposed. A mathematical discussion on the relationship between SIR-GCN and widely used GNNs is then laid out to put the contribution into context, establishing SIR-GCN as a generalization of classical GNN methodologies. Experiments on synthetic and benchmark datasets then demonstrate the relative superiority of SIR-GCN, outperforming comparable models in node and graph property prediction tasks.
@@ 1 Introduction
Graph neural networks (GNNs) constitute a class of deep learning models designed to process data that may be represented as graphs. These models are well-suited for node, edge, and graph property prediction tasks across various domains including social networks, molecular graphs, and biological networks, among others (Dwivedi et al., 2023; Hu et al., 2020). GNNs predominantly follow the message-passing scheme wherein each node aggregates the feature representation of its neighbors and combines them to create an updated node feature representation (Gilmer et al., 2017; Xu et al., 2018a; b). This allows the model to encapsulate both the network structure and the broader node contexts. Moreover, a graph readout function is employed to pool the individual node feature representation and create a representation for the entire graph (Li et al., 2015; Murphy et al., 2019; Xu et al., 2018a; Ying et al., 2018).
@@
Among the most widely used GNNs in literature include the graph convolution network (GCN) (Kipf & Welling, 2016), graph sample and aggregate (GraphSAGE) (Hamilton et al., 2017), graph attention network (GAT) (Brody et al., 2021; Veličković et al., 2017), and graph isomorphism network (GIN) (Xu et al., 2018a) which largely fall under the message-passing neural network (MPNN) (Gilmer et al., 2017) framework. These models have gained popularity due to their simplicity and remarkable performance across various applications (Dwivedi et al., 2023; Hsu et al., 2021; Hu et al., 2020; Jiang et al., 2022; Kim & Ye, 2020; Liu et al., 2020). Improvements are also constantly being proposed to achieve state-of-the-art performance (Bodnar et al., 2021; Bouritsas et al., 2022; Ishiguro et al., 2019; Miao et al., 2022; Sun et al., 2020; Wang et al., 2019b; Ying et al., 2021).
@@
Notably, these advances are mainly driven by heuristics and empirical results. Nonetheless, several studies have also begun exploring the representational capability of GNNs (Azizian & Lelarge, 2020; Bodnar et al., 2021; Böker et al., 2024; Corso et al., 2020; Garg et al., 2020; Sato et al., 2021). Most of these works analyzed GNNs in relation to the graph isomorphism task. Xu et al. (2018a) was among the first to lay the foundations for creating a maximally expressive GNN based on the Weisfeiler-Leman (WL) graph isomorphism test (Weisfeiler & Leman, 1968). Subsequent works build upon their results by considering extensions to the original 1-WL test. However, these results only hold with countable node feature representation which potentially limits their applicability. Meanwhile, Corso et al. (2020) proposed using multiple aggregators to create powerful GNNs when the space of node feature representation is uncountable. Interestingly, there has been no significant theoretical progress since this work.
@@
This paper presents a novel perspective on the representational capability of GNNs when the space of node feature representation is uncountable. The key idea is to define a pseudometric distance on the space of input to create a soft-injective function such that distinct inputs may produce similar outputs if and only if the distance between the inputs is sufficiently small on some representation. This idea is investigated across all levels—node-level, neighborhood-level, and graph-level. From the theoretical results, a simple and computationally efficient soft-isomorphic relational graph convolution network (SIR-GCN) which emphasizes the contextualized transformation of neighborhood feature representations using anisotropic and dynamic message functions is proposed. The mathematical relationship between SIR-GCN and popular GNNs in literature is also presented to underscore the novelty and advantages of the proposed model. Experiments on synthetic and benchmark datasets in node and graph property prediction tasks then highlight the expressivity of SIR-GCN, positioning the proposed model as the best-performing MPNN instance.
@@ 2 Graph neural networks
Let $\mathcal{G}=\left(\mathcal{V}_{\mathcal{G}},\mathcal{E}_{\mathcal{G}}\right)$ be a graph and $\mathcal{N}_{\mathcal{G}}(u)\subseteq\mathcal{V}_{\mathcal{G}}$ the set of nodes adjacent to node $u\in\mathcal{V}_{\mathcal{G}}$. The subscript $\mathcal{G}$ will be omitted whenever the context is clear. Suppose $\mathcal{H}$ is the space of node feature representation, henceforth feature, and $\bm{h_{u}}\in\mathcal{H}$ is the feature of node $u$. A GNN following the message-passing scheme can be expressed mathematically as
@@
where AGG and COMB are some aggregation and combination strategies, respectively, $\bm{H_{u}}$ is the multiset (Xu et al., 2018a) of neighborhood features for node $u$, $\bm{a_{u}}$ is the aggregated neighborhood features for node $u$, and $\bm{h^{*}_{u}}$ is the updated feature for node $u$. Since AGG takes arbitrary-sized multisets of neighborhood features as input and transforms them into a single feature, it may be considered a hash function. Hence, aggregation and hash functions shall be used interchangeably throughout the paper.
@@
When $\mathcal{H}$ is countable, Xu et al. (2018a) showed that there exists a function $f:\mathcal{H}\rightarrow\mathcal{S}$ such that the aggregation or hash function
@@
is injective or unique for each multiset of neighborhood features $\bm{H}$ of bounded size in the embedded feature space $\mathcal{S}$. This result forms the theoretical basis of GIN.
Meanwhile, the result above no longer holds when $\mathcal{H}$ is uncountable. In this setting, Corso et al. (2020) proved that if $\bigoplus$ comprises multiple aggregators (e.g., mean, standard deviation, max, and min), the hash function
@@
produces a unique output for every $\bm{H}$ of bounded size. This finding provides the foundation for the principal neighborhood aggregation (PNA) (Corso et al., 2020). Notably, for this result to hold, the number of aggregators in $\bigoplus$ must also scale with the size of the multiset of neighborhood features $\bm{H}$, which may be infeasible for large and dense graphs.
@@ 3 Soft-injective functions
While injective functions and metrics are necessary for tasks requiring strict isomorphism, many practical applications of GNNs often do not require such strict constraints. For instance, in node classification tasks, the model must produce identical outputs for some distinct nodes. Thus, this paper softly relaxes these constraints by employing pseudometrics and soft-injective functions.
Definition 1 (Pseudometric).
Let $\mathcal{H}$ be a non-empty set. A function $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ is a pseudometric on $\mathcal{H}$ if the following holds for all $\bm{h}^{(1)},\bm{h}^{(2)},\bm{h}^{(3)}\in\mathcal{H}$:
  • $d\left(\bm{h}^{(1)},\bm{h}^{(1)}\right)=0$;
[…]
Assumption 1.
Let $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ be a pseudometric on $\mathcal{H}$ such that $-d^{2}$ is a conditionally positive definite kernel on $\mathcal{H}$.
The Euclidean distance is an example of a pseudometric satisfying Assumption 1. A class of pseudometrics satisfying this assumption is provided below, see Berg et al. (1984) and Schölkopf (2000) for more.
Remark 1.
Consider the pseudometrics $d_{1}$ and $d_{2}$ on $\mathcal{H}$ satisfying Assumption 1. For $a>0$ and $10<p<$ 1, the pseudometrics $a\cdot d_{1}$, $\sqrt{d_{1}^{2}+d_{2}^{2}}$, and $d_{1}^{p}$ also satisfy Assumption 1.
Assumption 1 thus offers considerable flexibility in the choice of pseudometric $d$. The following theorem then softly relaxes the injective and metric requirements in previous works.
@@
Let $\mathcal{H}$ be a non-empty set with a pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{h}^{(1)},\bm{h}^{(2)}\in\mathcal{H}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
$\varepsilon_{2}<\left\lVert g\left(\bm{h}^{(1)}\right)-g\left(\bm{h}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<d\left(\bm{h}^{(1)},\bm{h}^{(2)}\right)<\varepsilon_{1}.$ (4)
@@
(a) Input feature space $\mathcal{H}$.
$\varepsilon$$g_{u}\left({\color[rgb]{0.1171875,0.55078125,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.1171875,0.55078125,1}\bm{h_{u}}^{(1)}}\right)$$g_{u}\left({\color[rgb]{0.15625,0.51171875,0.8828125}\definecolor[named]{pgfstrokecolor}{rgb}{0.15625,0.51171875,0.8828125}\bm{h_{u}}^{(2)}}\right)$$g_{u}\left({\color[rgb]{0.2734375,0.390625,1}\definecolor[named]{pgfstrokecolor}{rgb}{0.2734375,0.390625,1}\bm{h_{u}}^{(3)}}\right)$
(b) Embedded feature space $\mathcal{S}$.
Figure 1: Pseudometric $d_{u}$ and the corresponding feature map $g_{u}$.
Theorem 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $d_{u}$ that represents a dissimilarity function on $\mathcal{H}$, possibly encoded with prior knowledge, there exists a corresponding feature map $g_{u}$ that maps distinct inputs $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}\in\mathcal{H}$ close in the embedded feature space $\mathcal{S}$ if and only if $d_{u}$ determines $\bm{h_{u}}^{(1)},\bm{h_{u}}^{(2)}$ to be sufficiently similar on some representation. The lower bound $\varepsilon_{2}$ asserts the ability of $g_{u}$ to separate elements of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while the upper bound $\varepsilon_{1}$ ensures $g_{u}$ maintains the relationship between elements of $\mathcal{H}$ with respect to $d_{u}$. The feature map $g_{u}$ may then be described as soft-injective. [^1] The pseudometric $d$ induces the equivalence class $\left[\bm{h}\right]_{d}\coloneqq\left\{\bm{h^{\prime}}\in\mathcal{H}:d\left(\bm{h},\bm{h^{\prime}}\right)=0\right\}$ with the quotient space $\mathcal{H}_{d}\coloneqq\mathcal{H}\setminus d=\left\{\left[\bm{h}\right]_{d}:\bm{h}\in\mathcal{H}\right\}$ such that $d$ becomes metric and the corresponding feature map $g$ becomes injective on $\mathcal{H}_{d}$ (Schoenberg, 1938). Hence, $g$ may be described as soft-injective. Corollary 1 extends this result for multisets.
3.1 Soft-isomorphic relational graph convolution network
@@
$D^{2}\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(1)}\\ \bm{h^{\prime}}\in\bm{H}^{(1)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})-\dfrac{1}{2}\sum_{\begin{subarray}{c}\bm{h}\in\bm{H}^{(2)}\\ \bm{h^{\prime}}\in\bm{H}^{(2)}\end{subarray}}d^{2}(\bm{h},\bm{h^{\prime}})$ (5)
for some pseudometric $d:\mathcal{H}\times\mathcal{H}\rightarrow\mathbb{R}_{\geq 0}$ satisfying Assumption 1 and bounded, equinumerous multisets $\bm{H}^{(1)},\bm{H}^{(2)}$. There exists a feature map $g:\mathcal{H}\rightarrow\mathcal{S}$ such that for every $\bm{H}^{(1)},\bm{H}^{(2)}$ and $\varepsilon_{1}>\varepsilon_{2}>0$,
$\varepsilon_{2}<\left\lVert G\left(\bm{H}^{(1)}\right)-G\left(\bm{H}^{(2)}\right)\right\rVert<\varepsilon_{1}\iff\varepsilon_{2}<D\left(\bm{H}^{(1)},\bm{H}^{(2)}\right)<\varepsilon_{1},$ (6)
@@
Similarly, Corollary 1 shows that, for each node $u\in\mathcal{V}$, given a pseudometric distance $D_{u}$ on multisets of $\mathcal{H}$ defined in Eqn. 5 with a corresponding pseudometric distance $d_{u}$ on $\mathcal{H}$, there exists a corresponding feature map $g_{u}$ and soft-injective hash function $G_{u}$ defined in Eqn. 7 that produces similar outputs for distinct multisets of neighborhood features $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ if and only if $D_{u}$ deems $\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)}$ to be sufficiently similar on some representation. Likewise, the lower and upper bounds guarantee the ability of $G_{u}$ to separate equinumerous multisets of $\mathcal{H}$ in the embedded feature space $\mathcal{S}$ while maintaining the relationship with respect to $D_{u}$. In this setting, the feature map $g_{u}$ may be interpreted as the message function (Gilmer et al., 2017) of the aggregation strategy that transforms the individual neighborhood features. Meanwhile, the psuedometric $D_{u}$ corresponds to the kernel distance (Joshi et al., 2011) which intuitively represents the difference between the cross-distance and self-distance between two multisets.
Dynamic transformation
To illustrate the role of pseudometrics, consider anomaly detection on node features representing zero-mean scores, with node $u$ having two neighbors $v_{1}$ and $v_{2}$. If $d_{u}$ simply corresponds to the Euclidean distance, then the corresponding hash function $G_{u}$ becomes linear as presented in Fig. 2(a). The contour plot highlights collisions—instances where distinct inputs produce identical outputs (i.e., the equivalence class $\left[\bm{H}\right]_{D}$)—between dissimilar multisets of neighborhood features, resulting in aggregated neighborhood features that are less useful for the task.
@@
Nevertheless, other choices of pseudometrics, possibly incorporating prior knowledge, would correspond to more complex message functions $g_{u}$. This leads to non-trivial hash functions $G_{u}$ and contour plots where only the regions determined by $D_{u}$ to be similar may produce similar aggregated neighborhood features, making collisions more informative and controlled. This also highlights the significance of dynamic (Brody et al., 2021) or non-linear message functions $g_{u}$ in MPNNs.
As further illustration, if $d_{u}$ instead corresponds to the Euclidean distance of the squared score, then the corresponding hash function $G_{u}$ has the contour plot in Fig. 2(b). The resulting hash collisions and equivalence classes then become more useful and meaningful for detecting anomalous scores.
Anisotropic messages
@@
It is also worth noting that Corollary 1 holds for each node $u\in\mathcal{V}$ independently. Hence, different nodes may correspond to different $D_{u}$, $d_{u}$, $g_{u}$, and $G_{u}$. For simplicity, especially in inductive learning contexts, consider a single pseudometric instead, defined as
$D^{2}\left(\bm{H_{u}}^{(1)},\bm{H_{u}}^{(2)};\bm{h_{u}}\right)=\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(1)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(1)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}})-\dfrac{1}{2}\!\!\sum_{\begin{subarray}{c}\bm{h}\in\bm{H_{u}}^{(2)}\\ \bm{h^{\prime}}\in\bm{H_{u}}^{(2)}\end{subarray}}\!\!d^{2}(\bm{h},\bm{h^{\prime}};\bm{h_{u}}),$ (8)
with a single hash function, defined as
$G\left(\bm{H_{u}};\bm{h_{u}}\right)=\sum_{\bm{h}\in\bm{H_{u}}}g\left(\bm{h};\bm{h_{u}}\right),$ (9)
for every node $u\in\mathcal{V}$. This approach makes $D$, $d$, $g$, and $G$ anisotropic (Dwivedi et al., 2023) (i.e., a function of both the features of the query (center) node $\bm{h_{u}}$ and key (neighboring) nodes $\bm{h}\in\bm{H_{u}}$). Moreover, contextualized on the features of the query node, $D$ may still be interpreted as a pseudometric controlling hash collisions with a corresponding soft-injective hash function $G$.
@@
Furthermore, the integration of $\bm{h_{u}}$ also allows for the interpretation of $g$ as a relational message function, guiding how features of the key nodes are to be embedded and transformed based on the features of the query node. Figs. 2(c) and 2(d) provide intuition for this idea where the introduction of a bias term, assuming a function of the features of the query node, shifts the contour plot to produce distinct aggregated neighborhood features $\bm{a_{u}}\neq\bm{a_{u^{\prime}}}$ for nodes $u$ and $u^{\prime}$ with identical neighborhood features $\bm{H_{u}}=\bm{H_{u^{\prime}}}$ but distinct features $\bm{h_{u}}\neq\bm{h_{u^{\prime}}}$. Nevertheless, one may also inject stochasticity into the node features to distinguish between nodes $u$ and $u^{\prime}$ with identical features $\bm{h_{u}}=\bm{h_{u^{\prime}}}$ and neighborhood features $\bm{H_{u}}=\bm{H_{u^{\prime}}}$ with high probability (Sato et al., 2021) and to imitate having distinct $D_{u}$, $d_{u}$, $g_{u}$, and $G_{u}$ for each node $u\in\mathcal{V}$.
Proposed model